OpenStudy (anonymous):

The perimeter of a rectangular lot is 32 meters. The length exceeds the width by 4 meters. Find the length of the lot. Answer 6 m 14 m 10 m 18 m

OpenStudy (anonymous):

P=2l+2w

OpenStudy (anonymous):

l=w+4 because it exceeds the width by 4 aka w+4

OpenStudy (anonymous):

I got 14.

OpenStudy (anonymous):

its actually 6

OpenStudy (anonymous):

ill set it all up one sec

OpenStudy (anonymous):

\[32=2(w+4)+2w \] \[32=2w+8+2w\] \[32=4w+8\] \[24=4w\] \[w=6\]

OpenStudy (anonymous):

you know that the length=w+4 that is why you plug it in for l

OpenStudy (anonymous):

@ramirokool makes sense?

OpenStudy (anonymous):

Yes